### v3

@@ 1 Introduction
Despite this, GpI seems quite a bit easier than GI. For example, Tarjan’s $n^{\log n+O(1)}$ algorithm for groups [Mil78] can now be given as an exercise to undergraduates: every group is generated by at most $\lceil\log_{2}|G|\rceil$ elements, so the algorithm is to try all possible $\binom{n}{\log n}\leq n^{\log n}$ generating sets, and for each, check in $n^{O(1)}$ time whether the map of generating sets extends to an isomorphism. In contrast, the quasi-polynomial time algorithm for graphs was a tour de force that built on decades of cutting-edge research into algorithms and the structure of permutation groups. Nonetheless, it remains unknown whether the problem for groups is actually easier than that for graphs, or even whether both problems are in $\mathsf{P}$!
Using a finer notion of reduction, Chattopadhyay, Torán, and Wagner [CTW13] proved that there was no $\mathsf{AC^{0}}$ reduction from GI to GpI. This gave the first (and still only known) unconditional evidence that there is some formal sense (namely, the $\mathsf{AC^{0}}$ sense) in which GpI really is easier than GI. The key to their result was that the generator-enumeration technique described above can be implemented by non-deterministically guessing $\log^{2}n$ bits (describing the $\log n$ generators, each of $\log n$ bits), and then verifying an isomorphism by a non-deterministic circuit of depth only $O(\log\log n)$, a class we denote $\exists^{\log^{2}n}\mathsf{FOLL}$. Observe that $\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\textsf{quasiFOLL}$ (by trying all $2^{O(\log^{2}n)}$ settings of the non-deterministic bits in parallel), which cannot compute Parity [Raz93, Smo87, CTW13]. As GI is $\mathsf{DET}$-hard [Tor04]—and hence can compute Parity—there can be no $\mathsf{AC^{0}}$ reduction from GI to GpI.
@@
In this paper, we aim to close the gap between $\mathsf{AC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}$ in the complexity of Group Isomorphism and related problems. Our goal is to obtain constant-depth circuits of quasipolynomial size, a natural benchmark in circuit complexity [Bar92]. We leave it for future work to focus on reducing the size of the circuits. Our first main result along these lines is:
@@
The prior best results in terms of depth complexity all had super-constant depth (with quasipolynomial size): $\exists^{\log^{2}n}\mathsf{SC}^{2}$ by Tang [Tan13] yields (by a standard simulation argument) circuits of depth $\log^{2}n$, while $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ [CTW13] has depth $\log\log n$.
@@
Another very natural problem in computational algebra is the Min Generating Set (MGS) problem. Given a group, this problem asks to find a generating set of the smallest possible size. Given that many algorithms on groups depend on the size of a generating set, finding a minimum generating set has the potential to be a widely applicable subroutine. Despite this, the MGS problem for groups was shown to be in $\mathsf{P}$ by Lucchini & Thakkar very recently [LT24]. We improve their complexity bound to:
@@
Papadimitriou and Yannakakis did not specify the type of reduction used in their conjecture, though their $\exists^{\log^{2}n}\mathsf{P}$-completeness result for Log Generating Set of a Magma works in both logspace and $\mathsf{AC^{0}}$ (under a suitable input encoding). Our two upper bounds rule out (unconditionally in one case, conditionally in the other) such reductions for MGS for Quasigroups:
@@
In strong contrast, we show that MGS for Magmas is $\mathsf{NP}$-complete (Thm. 7.13).
@@
Membership for quasigroups is in $\mathsf{NTIME}(\mathrm{polylog}(n))\subseteq\mathsf{quasiAC^{0}}$.
@@
  • By known $\mathsf{AC^{0}}$ reductions (see, e.g., Levet [Lev23] for details), our $\mathsf{quasiAC^{0}}$ analysis of Chattopadhyay, Torán, and Wagner’s algorithm for Quasigroup Isomorphism yields the same upper bound for the isomorphism problems for Steiner triple systems, pseudo-STS graphs, Latin square graphs, and Steiner $(t,t+1)$-designs, as well as Latin Square Isotopy. Prior to our work, Quasigroup Isomorphism was not known to be solvable using quasiAC circuits of depth $o(\log\log n)$.
  • Group Isomorphism for simple groups or for groups from a dense set $\Upsilon$ of orders can be solved in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}\cap\mathsf{L}\cap\mathsf{FOLL}$. For groups in a dense set of orders, this improves the parallel complexity compared to the original result of Dietrich & Wilson [DW22]. As in their paper, note that $\Upsilon$ omits large prime powers. Thus, we essentially have that for groups that are not $p$-groups, Group Isomorphism belongs to a proper subclass of DET. This evidence fits with the widely-believed idea that $p$-groups are a bottleneck case for Group Isomorphism.
  • Abelian Group Isomorphism is in $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. The key novelties here are (1) a new observation that allows us to reduce the number of co-nondeterministic bits from $\log n$ (as in [GL23]) down to $\log\log n$, and (2) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding, rather than $\mathsf{FOLL}$ as in [CTW13].
  • Membership for nilpotent groups is in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$.
@@ 1.1 Methods
One such instance is in our improved bound for order-finding and exponentiation in a semigroup (Lem. 3.1). The previous proof [BKLM01] (still state of the art 23 years later) used a then-novel and clever “double-barrelled” recursive approach to compute these in $\mathsf{FOLL}$. In contrast, our proof uses standard repeated doubling, noting that it can be done in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$, recovering their result with standard tools and reducing the depth. We use this improved bound on order-finding to improve the complexity of isomorphism testing of Abelian groups (Thm. 5.1), simple groups (Cor. 3.3), and groups of almost all orders (Thm. 6.1).
For a few of our results, however, we need to develop new tools to work with quasigroups. In particular, for the $\mathsf{quasiAC^{0}}$ upper bound on MGS for quasigroups, we cannot directly adapt the technique of Chattopadhyay, Torán, and Wagner. Indeed, their analysis of their algorithm already seems tight to us in terms of having depth $\Theta(\log\log n)$. The first key is Thm. D, putting Membership for quasigroups into $\mathsf{NTIME}(\mathrm{polylog}(n))$. To do this, we replace the use of a cube generating sequences from [CTW13] with something that is nearly as good for the purposes of MGS: we extend the Babai–Szemerédi Reachability Lemma [BS84, Thm. 3.1] from groups (its original setting) to get straight-line programs in quasigroups. Division in quasigroups is somewhat nuanced, e.g., despite the fact that for any $a,b$, there exists a unique $x$ such that $ax=b$, this does not necessarily mean that there is an element “$a^{-1}$” such that $x=a^{-1}b$, because of the lack of associativity. Our proof is thus a careful adaptation of the technique of Babai & Szemerédi, with a few quasigroup twists that result in a slightly worse, but still sufficient, bound.
@@ 1.2 Prior work
The best known runtime bound for GpI is $n^{(1/4)\log_{p}(n)+O(1)}$-time due to Rosembaum [Ros13] (see [LGR16, Sec. 2.2]), though this tells us little about parallel complexity. In addition to Tarjan’s result mentioned above [Mil78], Lipton, Snyder, & Zalcstein [LSZ77] independently observed that if a group is $d$-generated, then we can decide isomorphism by considering all possible $d$-element subsets. This is the generator enumeration procedure. Using the fact that every group admits a generating set of size $\leq\log_{p}(n)$ (where $p$ is the smallest prime dividing $n$), Tarjan obtained a bound of $n^{\log_{p}(n)+O(1)}$-time for Group Isomorphism, while Lipton, Snyder, & Zalcstein [LSZ77] obtained a stronger bound of $\textsf{DSPACE}(\log^{2}n)$. Miller [Mil78] extended Tarjan’s observation to the setting of quasigroups. There has been subsequent work on improving the parallel complexity of generator enumeration for quasigroups, resulting in bounds of $\exists^{\log^{2}n}\mathsf{AC}^{1}$ ($\mathsf{AC}^{1}$ circuits that additionally receive $O(\log^{2}n)$ non-deterministic bits, denoted by other authors as $\beta_{2}\mathsf{AC}^{1}$) due to Wolf [^7] Wolf actually claims a bound of $\exists^{\log^{2}n}\textsf{NC}^{2}$; however, he uses $\textsf{NC}^{1}$ circuits to multiply two elements of a quasigroup rather than $\mathsf{AC^{0}}$ circuits. [Wol94], $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ due to Wagner [Wag10], and $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ due to Chattopadhyay, Torán, & Wagner [CTW13]. In the special case of groups, generator enumeration is also known to belong to $\exists^{\log^{2}n}\textsf{SC}^{2}$ [Tan13]. There has been considerable work on polynomial-time, isomorphism tests for several families of groups, as well as more recent work on NC isomorphism tests—we refer to recent works [GQ17, DW22, GL23] for a survey. We are not aware of work on isomorphism testing for specific families of quasigroups that are not groups.
@@ 2.1 Algebra and Combinatorics
Quasigroups. A quasigroup consists of a set $G$ and a binary operation $\star:G\times G\to G$ satisfying the following. For every $a,b\in G$, there exist unique $x,y$ such that $a\star x=b$ and $y\star a=b$. We write $x=a\backslash b$ and $y=b/a$, i. e., $a\star(a\backslash b)=b$ and $(b/a)\star a=b$. When the multiplication operation is understood, we simply write $ax$ for $a\star x$. A sub-quasigroup of a quasigroup is a subset that itself is a quasigroup. This means it is closed under the multiplication as well as under left and right quotients. Given $X\subseteq G$, the sub-quasigroup generated by $X$ is denoted as $\langle X\rangle$. It is the smallest sub-quasigroup containing $X$.
@@
As quasigroups are non-associative, the parenthesization of a given expression may impact the resulting value. For a sequence $S:=(s_{0},s_{1},\ldots,s_{k})$ and parenthesization $P$ from a quasigroup, define:
@@
A Latin square of order $n$ is an $n\times n$ matrix $L$ where each cell $L_{ij}\in[n]$, and each element of $[n]$ appears exactly once in a given row or a given column. Latin squares are precisely the Cayley tables corresponding to quasigroups. We will abuse notation by referring to a quasigroup and its multiplication table interchangeably. An isotopy of Latin squares $L_{1}$ and $L_{2}$ is an ordered triple $(\alpha,\beta,\gamma)$, where $\alpha,\beta,\gamma:L_{1}\to L_{2}$ are bijections satisfying the following: whenever $ab=c$ in $L_{1}$, we have that $\alpha(a)\beta(b)=\gamma(c)$ in $L_{2}$. Alternatively, we may view $\alpha$ as a permutation of the rows of $L_{1}$, $\beta$ as a permutation of the rows of $L_{1}$, and $\gamma$ as a permutation of the values in the table. Here, $L_{1}$ and $L_{2}$ are isotopic precisely if (i) the $(i,j)$ entry of $L_{1}$ is the $(\alpha(i),\beta(j))$ entry of $L_{2}$, and (ii) $x$ is the $(i,j)$ entry of $L_{1}$ if and only if $\gamma(x)$ is the $(\alpha(i),\beta(j))$ entry of $L_{2}$.
For a given Latin square $L$ of order $n$, we associate a Latin square graph $G(L)$ that has $n^{2}$ vertices; one for each triple $(a,b,c)$ that satisfies $ab=c$. Two vertices $(a,b,c)$ and $(x,y,z)$ are adjacent in $G(L)$ precisely if $a=x$, $b=y$, or $c=z$. Miller showed that two Latin square graphs $G_{1}$ and $G_{2}$ are isomorphic if and only if the corresponding Latin squares, $L_{1}$ and $L_{2}$, are main class isotopic; that is, if $L_{1}$ and $L_{2}$ can be placed into compatible normal forms that correspond to isomorphic quasigroups [Mil78]. A Latin square graph on $n^{2}$ vertices is a strongly regular graph with parameters $(n^{2},3(n-1),n,6)$. Conversely, a strongly regular graph with these same parameters $(n^{2},3(n-1),n,6)$ is called a pseudo-Latin square graph. Bruck showed that for $n>23$, a pseudo-Latin square graph is a Latin square graph [Bru63].

### v4

@@ 1 Introduction
Despite this, GpI seems quite a bit easier than GI. For example, Tarjan’s $n^{\log n+O(1)}$ algorithm for groups [Mil78] can now be given as an exercise to undergraduates: every group is generated by at most $\lceil\log_{2}|G|\rceil$ elements, so the algorithm is to try all possible $\binom{n}{\log n}\leq n^{\log n}$ generating sets, and for each, check in $n^{O(1)}$ time whether the map of generating sets extends to an isomorphism. In contrast, the quasi-polynomial time algorithm for graphs was a tour de force that built on decades of research into algorithms and the structure of permutation groups. Nonetheless, it remains unknown whether the problem for groups is actually easier than that for graphs under polynomial-time reductions, or even whether both problems are in $\mathsf{P}$!
Using a finer notion of reduction, Chattopadhyay, Torán, and Wagner [CTW13] proved that there was no $\mathsf{AC^{0}}$ reduction from GI to GpI. This gave the first unconditional evidence that there is some formal sense (namely, the $\mathsf{AC^{0}}$ sense) in which GpI really is easier than GI. The key to their result was that the generator-enumeration technique described above can be implemented by non-deterministically guessing $\log^{2}n$ bits (describing the $\log n$ generators, each of $\log n$ bits), and then verifying an isomorphism by a non-deterministic circuit of depth only $O(\log\log n)$, a class we denote $\exists^{\log^{2}n}\mathsf{FOLL}$. Observe that $\exists^{\log^{2}n}\mathsf{FOLL}\subseteq\textsf{quasiFOLL}$ (by trying all $2^{O(\log^{2}n)}$ settings of the non-deterministic bits in parallel), which cannot compute Parity [Hås86], even if augmented with $MOD_{p}$ gates for $p$ an odd prime [Raz93, Smo87, CTW13]. As GI is $\mathsf{DET}$-hard [Tor04]—and hence can compute Parity—there can be no $\mathsf{AC^{0}}$ reduction from GI to GpI.
@@
In this paper, we aim to close the gap between $\mathsf{AC^{0}}$ and $\exists^{\log^{2}n}\mathsf{FOLL}$ in the complexity of Group Isomorphism and related problems. Our goal is to obtain constant-depth circuits of quasipolynomial size, a natural benchmark in circuit complexity [Bar92]. Significantly improving the size of these circuits would improve the state of the art run-time of Group Isomorphism, a long-standing open question that we do not address here.
Our first main result along these lines is:
@@
Focusing on depth bounds—that is, without attempting to improve the worst-case runtime—our result is close to the end of the line for a series of works stretching back to 1970; see Table 1. The only possible further improvements we see, without improving the worst-case runtime, are to get an $\exists^{\log^{2}n}\mathsf{AC^{0}}$ upper bound—for which there are many obstacles, see Section 8—or improving the exact size and depth of our result, but here there is not much room for improvement, as we already get quasi-polynomial-size circuits of depth only 4 and size $n^{O(\log^{c}n)}$ for some $1<c\leq 4$.
Year Result Depth Citation
1970 Generator-enumerator introduced $\operatorname{poly}(n)$ Felsch & Neubüser [FN70] [^4] Complexity not analyzed there, but the same as Tarjan’s algorithm [Mil78].
1978 $\exists^{\log^{2}n}\mathsf{P}\subseteq\mathsf{DTIME}(n^{\log n+O(1)})$ $\operatorname{poly}(n)$ Tarjan (see Miller [Mil78])
1977 $\mathsf{DSPACE}(\log^{2}n)$ $O(\log^{2}n)$ Lipton–Snyder–Zalcstein [LSZ77] [^5] Despite the publication dates, this seems to have been independent of Tarjan’s result. They note that Miller and Rabin had also observed this result independently.
1994 $\exists^{\log^{2}n}\mathsf{AC}^{1}$ $O(\log n)$ Wolf [Wol94] [^6] Wolf only claims a bound of $\exists^{\log^{2}n}\textsf{NC}^{2}$. However, if we replace his use of $\textsf{NC}^{1}$ circuits to multiply two elements of a quasigroup with $\mathsf{AC^{0}}$ circuits, we immediately get the $\exists^{\log^{2}n}\mathsf{AC}^{1}$ bound.
2010 $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ $O(\log n)$ Wagner [Wag10]
2010 $\exists^{\log^{2}n}\mathsf{FOLL}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log\log n)$ Chattopadhyay–Torán–Wagner [CTW13] [^7] They do not claim the $\exists^{\log^{2}n}\mathsf{L}$ bound, but it follows immediately from their algorithm and results.
2013 $\exists^{\log^{2}n}\mathsf{SC}^{2}\cap\exists^{\log^{2}n}\mathsf{L}$ $O(\log n)$ Papakonstantinou–Tang–Qiao [Tan13] [^8] We have written the result this way, despite $\exists^{\log^{2}n}\mathsf{L}\subseteq\exists^{\log^{2}n}\mathsf{SC}^{2}$, because in Tang’s thesis [Tan13], the only place this is currently published, they only claim $\mathsf{NSC}^{2}$ using only $O(\log^{2}n)$ bits of nondeterminism, which in our notation would be $\exists^{\log^{2}n}\mathsf{SC}^{2}$. However, their algorithm and results also immediately yields a $\exists^{\log^{2}n}\mathsf{L}$ bound.
2024 $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ $\subseteq\mathsf{quasiAC^{0}}$ 4 This work
Table 1: History of the low-level circuit complexity of algorithms for Group Isomorphism based on the generator-enumerator technique. For non-circuit classes, we list their depth as the best-known depth of their simulation by circuits. The class in our bound is contained in all the other classes listed in the table. Although depth 4 in our result does not follow from the complexity class as listed here, it follows from the more exact class we use in Theorem 6.1.
We note that although there are depth reduction techniques for bounded-depth circuits with mod-$p$ gates [AH94], no such result is known for $\mathsf{quasiAC^{0}}$ circuits (without mod-$p$ gates).
@@
Another very natural problem in computational algebra is the Min Generating Set (MGS) problem. Given a group, this problem asks to find a generating set of the smallest possible size. Given that many algorithms on groups depend on the size of a generating set, finding a minimum generating set has the potential to be a widely applicable subroutine. The MGS problem for groups was shown to be in $\mathsf{P}$ by Lucchini & Thakkar only very recently [LT24]. We improve their complexity bound to:
@@
Papadimitriou and Yannakakis did not specify the type of reduction used in their conjecture, though their $\exists^{\log^{2}n}\mathsf{P}$-completeness result for Log Generating Set of a Magma works in both logspace and $\mathsf{AC^{0}}$ (under a suitable input encoding). Our two upper bounds rule out such reductions for MGS for Quasigroups (unconditionally in one case, conditionally in the other):
@@
In strong contrast, we show that MGS for Magmas is $\mathsf{NP}$-complete (Thm. 7.12).
@@
Membership for quasigroups is in $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{quasiAC^{0}}$.
@@
  • By known $\mathsf{AC^{0}}$ reductions (see, e.g., Levet [Lev23] for details), our $\mathsf{quasiAC^{0}}$ analysis of Chattopadhyay, Torán, and Wagner’s algorithm for Quasigroup Isomorphism yields the same upper bound for the isomorphism problems for Steiner triple systems, pseudo-STS graphs, Latin square graphs, and Steiner $(t,t+1)$-designs, as well as Latin Square Isotopy. Prior to our work, Quasigroup Isomorphism was not known to be solvable using quasiAC circuits of depth $o(\log\log n)$. See Cor. 6.2.
  • Group Isomorphism for simple groups (Cor. 3.3) or for groups from a dense set $\Upsilon$ of orders (Thm. 5.1) can be solved in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}\cap\mathsf{L}\cap\mathsf{FOLL}$. For groups in a dense set of orders, this improves the parallel complexity compared to the original result of Dietrich & Wilson [DW22]. As in their paper, note that $\Upsilon$ omits large prime powers. Thus, we essentially have that for groups that are not $p$-groups, Group Isomorphism belongs to a proper subclass of DET. This evidence fits with the widely-believed idea that $p$-groups are a bottleneck case for Group Isomorphism.
  • Abelian Group Isomorphism (Thm. 4.1) is in $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. The key novelties here are (1) a new observation that allows us to reduce the number of co-nondeterministic bits from $\log n$ (as in [GL23]) down to $\log\log n$, and (2) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding, rather than $\mathsf{FOLL}$ as in [CTW13].
  • Membership for nilpotent groups is in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$ (Prop. 7.3).
@@ 1.1 Methods
One such instance is in our improved bound for order-finding and exponentiation in a semigroup (Lem. 3.1). The previous proof [BKLM01] (still state of the art 23 years later) used a then-novel and clever “double-barrelled” recursive approach to compute these in $\mathsf{FOLL}$. In contrast, our proof uses standard repeated doubling, noting that it can be done in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$, recovering their result with standard tools and reducing the depth; in fact, from our proof we get $\mathsf{quasiAC^{0}}$ circuits of depth 2, which is clearly optimal from the perspective of depth. We use this improved bound on order-finding to improve the complexity of isomorphism testing of Abelian groups (Thm. 4.1), simple groups (Cor. 3.3), and groups of almost all orders (Thm. 5.1).
For a few of our results, however, we need to develop new tools to work with quasigroups. In particular, for the $\mathsf{quasiAC^{0}}$ upper bound on MGS for quasigroups, we cannot directly adapt the technique of Chattopadhyay, Torán, and Wagner. Indeed, their analysis of their algorithm already seems tight to us in terms of having depth $\Theta(\log\log n)$.
The first key is Thm. D, putting Membership for quasigroups into $\mathsf{NTIME}(\mathrm{polylog}(n))$ (more precisely, into $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$). To do this, we replace the use of a cube generating sequences from [CTW13] with something that is nearly as good for the purposes of MGS: we extend the Babai–Szemerédi Reachability Lemma [BS84, Thm. 3.1] from groups (its original setting) to quasigroups in order to obtain what we call cube-like generating sequences, which also give us short straight-line programs. Division in quasigroups is somewhat nuanced, e.g., despite the fact that for any $a,b$, there exists a unique $x$ such that $ax=b$, this does not necessarily mean that there is an element “$a^{-1}$” such that $x=a^{-1}b$, because of the lack of associativity. Our proof is thus a careful adaptation of the technique of Babai & Szemerédi, with a few quasigroup twists that result in a slightly worse, but still sufficient, bound.
@@ 1.2 Prior work
The best known runtime bound for GpI is $n^{(1/4)\log_{p}(n)+O(1)}$-time due to Rosembaum [Ros13] (see [LGR16, Sec. 2.2]), though this tells us little about parallel complexity. In addition to Tarjan’s result mentioned above [Mil78], Lipton, Snyder, & Zalcstein [LSZ77] independently observed that if a group is $d$-generated, then we can decide isomorphism by considering all possible $d$-element subsets. This is the generator enumeration procedure. Using the fact that every group admits a generating set of size $\leq\log_{p}(n)$ (where $p$ is the smallest prime dividing $n$), Tarjan obtained a bound of $n^{\log_{p}(n)+O(1)}$-time for Group Isomorphism, while Lipton, Snyder, & Zalcstein [LSZ77] obtained a stronger bound of $\textsf{DSPACE}(\log^{2}n)$. Miller [Mil78] extended Tarjan’s observation to the setting of quasigroups. There has been subsequent work on improving the parallel complexity of generator enumeration for quasigroups, resulting in bounds of $\exists^{\log^{2}n}\mathsf{AC}^{1}$ ($\mathsf{AC}^{1}$ circuits that additionally receive $O(\log^{2}n)$ non-deterministic bits, denoted by other authors as $\beta_{2}\mathsf{AC}^{1}$) due to Wolf [Wol94], [^12] See footnote 6. $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ due to Wagner [Wag10], and $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ due to Chattopadhyay, Torán, & Wagner [CTW13]. In the special case of groups, generator enumeration is also known to belong to $\exists^{\log^{2}n}\textsf{SC}^{2}$ [Tan13]. There has been considerable work on polynomial-time, isomorphism tests for several families of groups, as well as more recent work on NC isomorphism tests—we refer to recent works [GQ17, DW22, GL23] for a survey. We are not aware of work on isomorphism testing for specific families of quasigroups that are not groups.
@@ 2.1 Algebra and Combinatorics
Quasigroups and Latin squares. A quasigroup consists of a set $G$ and a binary operation $\star:G\times G\to G$ satisfying the following. For every $a,b\in G$, there exist unique $x,y$ such that $a\star x=b$ and $y\star a=b$. We write $x=a\backslash b$ and $y=b/a$, i. e., $a\star(a\backslash b)=b$ and $(b/a)\star a=b$. When the multiplication operation is understood, we simply write $ax$ for $a\star x$. A sub-quasigroup of a quasigroup is a subset that itself is a quasigroup. This means it is closed under the multiplication as well as under left and right quotients. Given $X\subseteq G$, the sub-quasigroup generated by $X$ is denoted as $\langle X\rangle$. It is the smallest sub-quasigroup containing $X$.
@@
As quasigroups need not be associative, the parenthesization of a given expression may impact the resulting value. For a sequence $S:=(s_{0},s_{1},\ldots,s_{k})$ and parenthesization $P$ from a quasigroup, define:
@@
A Latin square of order $n$ is an $n\times n$ matrix $L$ where each cell $L_{ij}\in[n]$, and each element of $[n]$ appears exactly once in a given row or a given column. Latin squares are precisely the Cayley tables corresponding to quasigroups. We will abuse notation by referring to a quasigroup and its multiplication table interchangeably. An isotopy of Latin squares $L_{1}$ and $L_{2}$ is an ordered triple $(\alpha,\beta,\gamma)$, where $\alpha,\beta,\gamma:L_{1}\to L_{2}$ are bijections satisfying the following: whenever $ab=c$ in $L_{1}$, we have that $\alpha(a)\beta(b)=\gamma(c)$ in $L_{2}$. Alternatively, we may view $\alpha$ as a permutation of the rows of $L_{1}$, $\beta$ as a permutation of the columns of $L_{1}$, and $\gamma$ as a permutation of the values in the table. Here, $L_{1}$ and $L_{2}$ are isotopic precisely if $x$ is the $(i,j)$ entry of $L_{1}$ if and only if $\gamma(x)$ is the $(\alpha(i),\beta(j))$ entry of $L_{2}$.
@@
A Latin square $L$ can equivalently be viewed as a set of triples $\{(i,j,L_{ij}):i,j\in[n]\}\subseteq[n]\times[n]\times[n]$. Given a set of triples $S\subseteq[n]\times[n]\times[n]$, the Latin square property can equivalently be rephrased as: every $i\in[n]$ appears as the first—resp. second, resp. third—coordinate of some triple in $S$, and no two triples in $S$ agree in more than one coordinate. From this perspective, an additional potential symmetry of Latin squares emerges: two Latin squares $L_{1},L_{2}$ are parastrophic [^13] This terminology is borrowed from the quasigroup literature. In the Latin square literature this is sometimes referred to as “conjugate”, but we find the argument in Keedwell and Denes [KD15, pp. 15–16] to use the quasigroup nomenclature even in the setting of Latin squares compelling. if there is a permutation $\pi\in S_{3}$ such that, when viewed as sets of triples, we have $L_{2}=\{(x_{1^{\pi}},x_{2^{\pi}},x_{3^{\pi}}):(x_{1},x_{2},x_{3})\in L_{1}\}$; the induced map $L_{1}\to L_{2}$ is called a parastrophy. A main class isomorphism of Latin squares is the composition of a parastrophy and an isotopy; if there exists a main class isomorphism $L_{1}\to L_{2}$, we say they are main class isomorphic. [^14] Miller [Mil78], and then Levet [Lev23] following Miller, referred to this as “main class isotopy”; we have since found several modern textbooks on quasigroups and Latin squares that refer to this notion as one of “main class isomorphism”, “paratopy”, or “isostrophy” (sic). Keedwell and Denes [KD15, pp. 15–16] have a nice discussion of the terminology, as well as the history of its usage.
For a given Latin square $L$ of order $n$, we associate a Latin square graph $G(L)$ that has $n^{2}$ vertices; one for each triple $(a,b,c)$ that satisfies $ab=c$. Two vertices $(a,b,c)$ and $(x,y,z)$ are adjacent in $G(L)$ precisely if $a=x$ or $b=y$ or $c=z$. Miller showed that two Latin square graphs $G_{1}$ and $G_{2}$ are isomorphic if and only if the corresponding Latin squares, $L_{1}$ and $L_{2}$, are main class isomorphic [Mil78].
A Latin square graph on $n^{2}$ vertices is a strongly regular graph with parameters $(n^{2},3(n-1),n,6)$. Conversely, a strongly regular graph with these same parameters $(n^{2},3(n-1),n,6)$ is called a pseudo-Latin square graph. Bruck showed that for $n>23$, a pseudo-Latin square graph is a Latin square graph [Bru63].
@@
Designs. Let $t\leq k\leq v$ and $\lambda$ be positive integers. A $(t,k,\lambda,v)$ design is an incidence structure $\mathcal{D}=(X,\mathcal{B},I)$, where $X$ is a set of $v$ points, $\mathcal{B}$ is a subset of $\binom{X}{k}$—whose elements are referred to as blocks—and such that each $t$-element subset of $X$ belongs to exactly $\lambda$ blocks. Now $I$ is the point-block incidence matrix, where $I_{x,B}=1$ precisely if the point $x$ belongs to the block $B$.
If $t<k<v$, we say that the design is non-trivial. If $\lambda=1$, the design is referred to as a Steiner design. We denote Steiner designs as $(t,k,v)$-designs when we want to specify $v$ the number of points, or Steiner $(t,k)$-designs when referring to a family of designs. We note that Steiner $(2,3)$-designs are known as Steiner triple systems. Projective planes are Steiner $(2,q+1,q^{2}+q+1)$-designs, and affine planes are Steiner $(2,q,q^{2})$-designs. We assume that designs are given by the point-block incidence matrix.
For a design $\mathcal{D}=(X,\mathcal{B},I)$, we may define a block intersection graph (also known as a line graph) $G(V,E)$, where $V(G)=\mathcal{B}$ and two blocks $B_{1},B_{2}$ are adjacent in $G$ if and only if $B_{1}\cap B_{2}\neq\emptyset$. In the case of a Steiner $2$-design, the block-intersection graph is strongly regular. For Steiner triple systems, the block-intersection graphs are strongly regular with parameters $(n(n-1)/6,3(n-3)/2,(n+3)/2,9)$. Conversely, strongly regular graphs with the parameters $(n(n-1)/6,3(n-3)/2,(n+3)/2,9)$ are referred to as pseudo-STS graphs. Bose showed that pseudo-STS graphs graphs with strictly more than $67$ vertices are in fact STS graphs [Bos63].